OpenStudy (anonymous):

The length of the base of the triangle is b, the length of the corresponding altitude is h, and the area is A. Find the missing part. base is 2 square root of 3 height is 3 square root of 3 what is the area in. sq. in.

OpenStudy (anonymous):

area is the base times the height divided by two

OpenStudy (anonymous):

so it's \[(2\sqrt{3})\times(3\sqrt{3})/2\] = \[18/2 = 9\]

OpenStudy (anonymous):



OpenStudy (anonymous):

how did you get 18

OpenStudy (anonymous):

2x3 = 6 \[\sqrt{3}\times \sqrt{3} = 3\] 6x3 = 18

OpenStudy (anonymous):

okay so u have to break the quare root down too

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

if you had another problem your base is ____________ your height is 5 and your area is 10 square inches I would have thought your base would be 2 but it keeps saying its wrong do you know what i am doing wrong

OpenStudy (lgbasallote):

\[\large A = \frac{bh}{2}\] \[\large 2A = bh\] \[\large \frac{2A}{h} = b\]

OpenStudy (anonymous):

^what @igbasallote is saying is that you have to multply by two to take the half factor into account

OpenStudy (anonymous):

so your answer is 4

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

The lengths of the sides of a right triangle are a and b, and the hypotenuse is c. Find the area of the triangle. a = 6 in.; b = 4 in. A = sq. in I know there is a special formula for this one I cant remember which one it is , can you help with this